OpenStudy (anonymous):

7(r+t)=5r+59 i Need Work Done To

OpenStudy (anonymous):

\[7r+7t=5r+59\] \[7r-5r=59-7t\] \[2r=59-7t\] \[r=(59-7t)/2\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

yea whats the answer though

OpenStudy (anonymous):

r=(59-7t)/2

OpenStudy (anonymous):

thanks